No results Optionsv Cutnell, Physics, 11e sten Announcements Grace Period: 6 days left I Register Nov PRINTER VERSION 4 BACK NEXT Chapter 16, Problem 012 ass of a string is 9.3 x 103 kg, and it is stretched so that the tension in it is 240 N. A transverse wave traveling on this string has a frequency of 180 Hz and a wavelength of 0.50 m. What is the length of the string? Number the tolerance is +/-2% Click if you would like to Show Work for this question: Units Open Show Work By accessing this Question Assistance, you will learn while you earn points based on the Point Potential Policy set by your instructor Question Attempts: Unlimited SAVE FOR LATER Earn Maximum Points available only if you answer this question correctly in four attempts or less. 4:26 PM
Answer #1

Frequency, f = 180 Hz

Wavelength, lambda =0.50m

Velocity of the wave, v = f * lambda = ( 180*0.5) =90 m/s

Velocity of the wave in stretched string, v = sqrt (T/ mu)

Where T= tension of the string = 240 N

mu= mass per unit length of the string

mu=(T/ (v^(2))) =(240 / (90*90)) = 0.02963 kg/m

Total mass of the string= 9.3* 10^(-3) kg

Length of the string, l = ( 9.3* 10^(-3))/(0.02963) = 0.3139 m = 31.39 cm (ans)